OpenStudy (anonymous):

I would love some help. Thanks in advance. How many solutions does the equation have? 4(2x+3)=2(3x-4)

OpenStudy (kirbykirby):

\[4(2x+3)=2(3x-4)\\ 8x+12=6x-8\\ 2x=-20\\ x=-10 \]

OpenStudy (kirbykirby):

how many solutions do you count :)

OpenStudy (anonymous):

Only one solution?

OpenStudy (kirbykirby):

\(\checkmark\)

OpenStudy (anonymous):

Awesome thank you! Do you have time for more?

OpenStudy (kirbykirby):

I guess lol

OpenStudy (anonymous):

How many solutions does the equation have?

OpenStudy (kirbykirby):

You would have infinitely many solutions. On the right side, you have \[ \frac{1}{5}(5a+25)\], distribute the 1/5: \[\frac{1}{5}(5a)+\frac{1}{5}(25)=a+5\] The right side and left side are exactly equal,. They are essentially the same equation. In this case plugging in any value for \(a\) will work! (To see this, your equation reduced to \[ a+5 = a+5\] Plug in a=1: \(1+5 = 1+5\) plug in a=2: \(2+5 = 2+5\) etc.

OpenStudy (anonymous):

Thank you for explaining. I understand a little more now. Here's the next one if you want to answer it. How many solutions does the equation have?

OpenStudy (kirbykirby):

\(3(4x-3)-7x\\ =3(4x)-3(3)-7x\\ =12x-9-7x\\ =5x-9\) This is equal to the right side of your equation. Can you guess how many solutions there are ;) ?

OpenStudy (anonymous):

One?

OpenStudy (kirbykirby):

you have the same equation on both sides, \(5x-9=5x-9\), it's the same situation as in the previous problem

OpenStudy (anonymous):

Infinite?

OpenStudy (anonymous):

This is the last one i need help with. Which expressions would complete this equation so that it has one solution? 4(2x – 3) – 4x = Choose exactly two answers that are correct. A. 4x – 12 B. 3x – 4 C. 3(2x – 6) D. 4(x – 5)

OpenStudy (anonymous):

@kirbykirby?

OpenStudy (anonymous):

You there?